OpenStudy (anonymous):

Using u substitution, integrate S x(sqrt(x^2 + 1))dx

OpenStudy (shiraz14):

I'm guessing that your S is actually the integral sign (∫)

OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

can you help me walk through this?

OpenStudy (shiraz14):

Use the substitution x = tan(u).

OpenStudy (shiraz14):

So x^2 + 1 = tan^2u + 1 = sec^2u & dx = sec^2u. Now, sub. these values into the integral & evaluate it. Can you try this?

OpenStudy (anonymous):

wait a minute, trig? it's a square root

OpenStudy (shiraz14):

We can use substitution to change a function in an integral - this is one of the intricacies of calculus ...

OpenStudy (anonymous):

u = x^2 + 1 du = 2x dx 1/2 du = x dx this seems much easier

OpenStudy (shiraz14):

@cleetus779 : Oh yes, you can do this as well ... :p

OpenStudy (anonymous):

came up with sqrt(u^3)/3 is that right?

OpenStudy (shiraz14):

It should be sqrt(u^3)/4 + c (where c is an arb. const).

OpenStudy (shiraz14):

But that's not the final answer - you should re-substitute the u back as a function of x, since the original question was using x, not u. Can you express your final answer in terms of x?

OpenStudy (anonymous):

ok, thanks for the help

OpenStudy (anonymous):

double checked here http://www.wolframalpha.com/input/?i=integrate+x%28sqrt%28x%5E2+%2B+1%29%29